OpenStudy (anonymous):

Where do I begin in a problem like this: y^2+3yz-8z-4x=0? where I have to solve for 0

OpenStudy (ghazi):

there are three variables but one equation ...so how you gonna get answer for one variable

OpenStudy (anonymous):

@ghazi ?

OpenStudy (anonymous):

What do you mean by "solve" for 0? You can't solve for a number. You already know what it is. Solve means you find out the value of an unknown, or put in terms of an unknown.

OpenStudy (ghazi):

x, y and z are three variable but you've got just one equation....how you gonna get the answer for each variable...the only option is to get one in terms of other

OpenStudy (anonymous):

@aseeran oh im sorry i have to solve for 'z'

OpenStudy (ghazi):

then separate out z from the equation

OpenStudy (anonymous):

y^2+3yz-8z-4x=0 3yz-8z=y^2+4x z(3y-8)=(-y^2+4x) z=(-y^2+4x)/(3y-8)

OpenStudy (anonymous):

i got z=4x-y^2/3y-8

OpenStudy (ghazi):

\[y^2 +3yz-8z-4x=0\] then \[z(3y-8)=4x+y^2\]

OpenStudy (ghazi):

\[z=\frac{ 4x+y^2 }{ 3y-8 }\]

OpenStudy (anonymous):

@sarah70 yeah ur answer is also fine

OpenStudy (anonymous):

ok thanks guys

OpenStudy (ghazi):

:)